OpenStudy (anonymous):

The area of a triangle is 3.6 cm2. If the triangle has a base of 6 cm, what is the height? 0.6 cm 1.2 cm 12 cm 3 cm

OpenStudy (anonymous):

1.2cm

OpenStudy (katrinakaif):

Remember to show your work. (Take a look at the COC)

OpenStudy (anonymous):

3.6x2/6=1.2cm

jimthompson5910 (jim_thompson5910):

A = (bh)/2 3.6 = (6h)/2 3.6*2 = 6h 7.2 = 6h 7.2/6 = h 1.2 = h h = 1.2 So the height is 1.2 cm